3. Mai 2016: Dienstag


  • Nightly Update: Two (2) new issues.
  • Accounts: No errors in 51 new accounts (since Friday).
  • Issues:
    • Two new OSP folks with different last names
    • Sent email to OSP folks for verification.
    • Confirmed – set up LDAP to find matches tonight
  • Development:
    • Rethinking LookAside in OU’s…
      • What-if: Instead of loading them into LDAP, include yesterday’s failures in today’s inputs?


  • Save Groups: The usual failures, plus license1
  • Clones: Started.
  • Tapes:
    • Replaced cleaning cartridge.
    • Labeled 40 more tapes into Scratch.


Leave a Reply

You must be logged in to post a comment.